OpenStudy (anonymous):

The height of male students at PWSH is approximately normally distributed with a mean of 71 inches and standard deviation of 2.5 inches. a) What percent of the male students are shorter than 66 inches? b) Taller than 73.5 inches? c) Between 66 & 73.5 inches?

OpenStudy (anonymous):

your sexxxy

OpenStudy (anonymous):

ok ok i need the answer

OpenStudy (anonymous):

please

jimthompson5910 (jim_thompson5910):

do you have a TI 83 or 84 calculator?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

alright, then wolfram alpha is the next best thing

jimthompson5910 (jim_thompson5910):

for part a) type in normalcdf(-100,66,71,2.5) to get the answer 0.02275

jimthompson5910 (jim_thompson5910):

the normalcdf function finds the area under the normal curve in general, you type in normalcdf(a,b,mu,sigma) 'a' and 'b' are the endpoints of where you want to find the area mu is the mean sigma is the standard deviation

OpenStudy (anonymous):

there is some formula to solve it? cause i dont know what normalcdf is

jimthompson5910 (jim_thompson5910):

there is no simple formula to memorize or do this by hand so you either have to use tables or a calculator

OpenStudy (anonymous):

what table?

jimthompson5910 (jim_thompson5910):

sometimes the book will have probability tables in it (or normal distribution tables)

jimthompson5910 (jim_thompson5910):

basically it gets really complicated to do it by hand, so you need a calculator before calculators were invented, tables were the only way to do it (in a reasonable amount of time)

OpenStudy (anonymous):

is there something about empirical or chebyshev rules?

jimthompson5910 (jim_thompson5910):

oh right, not sure why I didn't think about the empirical rule (it only works in a limited number of cases), but it looks like it will work here

OpenStudy (anonymous):

i know the answers already, but i don't know the way to find them

jimthompson5910 (jim_thompson5910):

what is the empirical rule? do you remember?

OpenStudy (anonymous):

it is about standarts deviaton from the mean

OpenStudy (anonymous):

99.7, 95, 68

jimthompson5910 (jim_thompson5910):

yes, good, 68% of the distribution is 1 standard deviation from the mean

OpenStudy (anonymous):

1 st deviation- 68 2-95 3-99.7

jimthompson5910 (jim_thompson5910):

95% of the distribution is 2 std deviations from the mean

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

99.7% of the distribution is 3 std deviations from the mean

jimthompson5910 (jim_thompson5910):

in this case, mu = 71 and sigma = 2.5 71 - 2.5 = 68.5 71 + 2.5 = 73.5 So 68% of the distribution is between 68.5 and 73.5

jimthompson5910 (jim_thompson5910):

71 - 2*2.5 = 66 71 + 2*2.5 = 76 and 95% of the distribution is between 66 and 76

jimthompson5910 (jim_thompson5910):

what percentage is left over?

OpenStudy (anonymous):

yes, you r right, i know that, but how yo solve a problem?

jimthompson5910 (jim_thompson5910):

I'm getting there

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

so 95% of the distribution is between 66 and 76 this means 100 - 95 = 5% is left over, agreed?

OpenStudy (anonymous):

i can give you the answers

jimthompson5910 (jim_thompson5910):

this 5% is distributed between the two tails so each tail gets 5/2 = 2.5%

jimthompson5910 (jim_thompson5910):

this means that roughly 2.5% of the population is shorter than 66 inches

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

what tail?

jimthompson5910 (jim_thompson5910):

ok I'll try to draw a picture of the distribution

jimthompson5910 (jim_thompson5910):